Kingston Company uses the dollar-value LIFO method of computing inventory. An external price index is used to convert ending inventory to base year. The company began operations on January 1, 2018, with an inventory of $255,000. Year-end inventories at year-end costs and cost indexes for its one inventory pool were as follows:

Year Ended Ending Inventory Cost Index
December 31 at Year-End Costs (Relative to Base Year)
2018 $ 319,300 1.03
2019 406,560 1.12
2020 384,770 1.09
2021 372,750 1.05


Required:
Calculate inventory amounts at the end of each year. (Round intermediate calculations and final answers to the nearest whole dollars.)
  

Ending Inventory Layers Converted to Base Year Cost Inventory Layers Converted to Cost Inventory DVL Cost Inventory at Year- End Cost Year-End Cost Index Inventory Layers at Base Year Cost Inventory Layers Converted to Cost Inventory Year-End Date Layers at Base Cost - Index Year Cost 01/01/2018 Base Base 2018 Base 2018 2019 Base 2018 2019 Base 2018 2019 2021 12/31/2018 12/31/2019 12/31/2020 12/31/2021
Answer #1

Calculation of Inventory amounts at the end of each year:

Inventory layers converted to base year cost Inventory layers converted to cost Ending Inventory DVL cost
Date inventory at year end cost / Year end cost index = inventory layers at base year cost Year inventory layers at base year cost * Year end cost index = inventory layers converted to cost for
1/1/18 255000 / 1 = 255000 255000 * 1 = 255000
12/31/18 319300 / 1.03 = 310000 255000 * 1 = 255000
2018 55000 * 1.03 = 56650 311650
12/31/19 406560 / 1.12 = 363000 255000 * 1 = 255000
2018 55000 * 1.03 = 56650
2019 53000 * 1.12 = 59360 371010
12/31/20 384770 1.09 = 353000 255000 * 1 = 255000
2018 55000 * 1.03 = 56650
2019 43000 * 1.12 = 48160 359810
12/31/21 372750 1.05 = 355000 255000 * 1 = 255000
2018 55000 * 1.03 = 56650
2019 45000 * 1.12 = 50400 362050